OpenStudy (anonymous):

Solve the limit... L= lim [a- √(a^2-x^2) -x^2/4]/x^4 x→0 a>0

OpenStudy (caozeyuan):

Holy cow! what a mess you ve got there

OpenStudy (anonymous):

wht ?? U havnt understood it ??

OpenStudy (caozeyuan):

\[\lim_{a \rightarrow 0 x \rightarrow 0}\frac{ a-\sqrt{a ^{2}-x ^{2}}-\frac{ x^{2} }{ 4 } }{ x ^{4}}\]

OpenStudy (caozeyuan):

Did I undestand correctly

OpenStudy (caozeyuan):

I think this can cause a Human Calculator run out of battery LOL

OpenStudy (caozeyuan):

sorry, a great than 0

hartnn (hartnn):

why don't you just plug in x = 0 ?

OpenStudy (caozeyuan):

but bottom is 0 if you do that @hartnn

OpenStudy (gorv):

well it is 0/0 form if plug x=0and a=0

OpenStudy (gorv):

and we can apply l hopital to it